OpenStudy (anonymous):

Solve 7^(x – 2) = 12^x

sam (.sam.):

7^(x-2)=12^(x) ln(7^(x-2))=ln(12^(x)) xln(7)-2ln(7)=ln(12^(x)) xln(7)-2ln(7)-xln(12)=0 x(ln(7)-ln(12))=2ln(7) x=(2ln(7))/(ln(7)-ln(12)) x=3.8918

OpenStudy (anonymous):

sam can you please help me with one more question, i would appreciate it

OpenStudy (anonymous):

whoa

sam (.sam.):

@aminah.love18 ? Anything wrong?

OpenStudy (anonymous):

no im fine for rnow

sam (.sam.):

ok :)

OpenStudy (anonymous):

sam can you help me with this question Solve 3^(x – 4) = 7^(x + 9)